OpenStudy (anonymous):

Use the graph of automobile styles owned by employees of a certain company to answer the question. What is the ratio of SUV or pick-up truck owners to owners of vans? A. 7:15 B. 1:2 C. 2:1 D. 15:7 Please help I'm sending attatchment

OpenStudy (anonymous):

@iGreen @k_lynn

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

"the ratio of SUV or pick-up truck owners to owners of vans" so look at the pie chart and tell me the number of employees that owns pick up truck and vans...

OpenStudy (anonymous):

80 if u add them

OpenStudy (anonymous):

no no, don't add them

OpenStudy (anonymous):

Ok 35 vans and 45 pick up trucks

OpenStudy (igreen):

SUVs or Pick-up Trucks 30 + 45 = 75 Vans 35 So the ratio is 75:35, can you simplify that?

OpenStudy (anonymous):

^ lol that explains it

OpenStudy (anonymous):

I don't think so

OpenStudy (igreen):

Just divide these: 75 / 5 35 / 5

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

75/5=15 35/5=7

OpenStudy (igreen):

Yes, so we have 15/7

OpenStudy (igreen):

That's our answer :)

OpenStudy (anonymous):

there is 35 vans not 30 when u said in ur problem

OpenStudy (anonymous):

or did we have to do that

OpenStudy (igreen):

Those are SUVs that was 30..I said vans were 35.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Thanks